OpenStudy (anonymous):

k=1sigma k tends to infinity (1/(k+1)(k+3)) how can i find this series is converges

OpenStudy (anonymous):

If the nth term tends to zero as n tends to infinity it converges

OpenStudy (anonymous):

plz solve the steps.?

OpenStudy (anonymous):

Wts the nth term?

OpenStudy (anonymous):

i dont know. how to find nth term of this sequence

OpenStudy (anonymous):

just replace k by n

OpenStudy (anonymous):

and then put n tending to infinity

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}1/(n+1)(n+3)\]

OpenStudy (zzr0ck3r):

if the nth term tends to 0 the divergence test is inconclusive

OpenStudy (anonymous):

put the limit then bottom value is infiny and 1/infinity is zero..? so can i find converges

OpenStudy (anonymous):

are u sure,...im rusty on this topic

OpenStudy (zzr0ck3r):

integral test, if you get a number out of the unbound integral it is divergenct and if not then its not

OpenStudy (anonymous):

see i dunno that much

OpenStudy (anonymous):

zzrocker plz can help me in this question..?

OpenStudy (zzr0ck3r):

can you do the integral of f(x) from 0 to infinity?

OpenStudy (zzr0ck3r):

err 1 to infinity?

OpenStudy (anonymous):

k=1 to infinity

OpenStudy (anonymous):

hellow ru there/?

OpenStudy (zzr0ck3r):

take the integral from 1 to infinity yes

OpenStudy (anonymous):

can u solve it..?

OpenStudy (anonymous):

in my question what is the closed form of the nth partil sum.?

OpenStudy (zzr0ck3r):

Im not sure what that is asking you, i think sigma 1->n f(x)

OpenStudy (anonymous):

k=1 sigma k tends to infity

OpenStudy (anonymous):

my question is how can i find closed form of the nth partial sum and whehter the series in coverges.?

OpenStudy (anonymous):

hellow any one see my queston

OpenStudy (anonymous):

We can't give you the answer sherazzzzz, we can only help you solve it.

OpenStudy (anonymous):

ok dear little bit help me plz

OpenStudy (zzr0ck3r):

if you do the integral you will see the series does not converge, because the integral does not converge. and sigma 1->n f(x) is how you write it

OpenStudy (zzr0ck3r):

sigma 1->n f(n)

OpenStudy (anonymous):

my book ans says converges to 1/2

OpenStudy (anonymous):

but how..? i dont know..:(

OpenStudy (zzr0ck3r):

you know that it converges because the inttegral converges. do you understand that?

OpenStudy (anonymous):

yes but how to find closed form of the nth partial sum.. this is direcect converges not to find nth partila sum

OpenStudy (anonymous):

hi buddy u understand my problem??

OpenStudy (zzr0ck3r):

no, I cant read what you wrote.

OpenStudy (anonymous):

find the closes form of the nth partial sum and determin wheter the seriers in converges?

OpenStudy (zzr0ck3r):

Find the closed form of the nth partial sum and determine whether the series converges?

OpenStudy (zzr0ck3r):

Do you understand that I have answered you about "if the series converges" ?

OpenStudy (anonymous):

how to find a closed form for the nth partial sum?

OpenStudy (anonymous):

yes borther but problem is how to find closed form for the nth partila sum?

OpenStudy (zzr0ck3r):

I think that is just saying use sigma notation? to write the function series expansion in terms of n with sigma notation>

OpenStudy (anonymous):

yes bro

OpenStudy (zzr0ck3r):

ok so I have said this many times as well sigma from 1 to n of f(n)

OpenStudy (anonymous):

\[\sum_{k=1}^{\infty} 1/(k+1)(k+3)\]

OpenStudy (zzr0ck3r):

1/((K+1)(K+3)) but yes, if that is what they are asking you to do, then that is the answer:)

OpenStudy (anonymous):

yar converges to 1/2 how..? con u know it.?

OpenStudy (zzr0ck3r):

is (k+1)(k+3) both in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

um there are many ways to see what it converges to, what are you doing in class?

OpenStudy (zzr0ck3r):

are you doing geometric series?

OpenStudy (anonymous):

brother i study in virtual university..

OpenStudy (anonymous):

no ;(

OpenStudy (anonymous):

Hey guys, what about comparision test? \[\sum_{k}^{\infty} 1/(k+1)(k+3) = \sum_{k}^{\infty} 1/(k^2+4k+3)\] \[1/(k^2+4k+3) < 1/k^2 \] so comparing with an armonic series as 1/k^2, as the bigger converge the smaller converge too

OpenStudy (zzr0ck3r):

he knows it converges, but how to get the sum is the question?

OpenStudy (anonymous):

brother not to find comparison ttest just find closed from of nth partial sum

OpenStudy (zzr0ck3r):

are you hulk hogan?

OpenStudy (anonymous):

bro plz help me ..:( plz how to find closed form of the nth partial sum

OpenStudy (zzr0ck3r):

I dont know. if its not geometric i dont know how to do it.

OpenStudy (zzr0ck3r):

but your book says 1/2 and wolfram says 5/12

OpenStudy (anonymous):

yes bro its not a geometrc serires..:(

OpenStudy (anonymous):

telescopic series

OpenStudy (anonymous):

you can make 1/(k+1)(k+3)=(1/2)[(1/(k+1)) - (1/(k+3))]

OpenStudy (anonymous):

and now you can solve as a telescopic serie

OpenStudy (zzr0ck3r):

O yeah , I sort of remember those. rewrite it as two series and look for canceling pattern.

OpenStudy (anonymous):

yeah I've done the serie and the sum is 5/12 so this is the way

OpenStudy (anonymous):

may I write it?

OpenStudy (zzr0ck3r):

yeah same here, i was gonna scan it but please go ahead:)

OpenStudy (zzr0ck3r):

first expand then write as 1/2sum(1/(x+1)) - 1/2sum(1/(x+3)) = 1/2(1/2+1/3+1/4+1/5......) - 1/2(1/4+1/5.....) cancel like terms and get 1/2(1/2+1/3) = 5/12

OpenStudy (zzr0ck3r):

the book is wrong:)

OpenStudy (anonymous):

ok so as I said above you can make 1/(k+1)(k+3)=(1/2)[(1/(k+1)) - (1/(k+3))]; now if you make k=1 the firs term is 1/2-1/4 if you give k=2 1/3-1/5; 1/4-1/5 (k=3); ... 1/k - 1/(k+2) ; (k-1 term) 1/(k+1) - 1/(k+3) ; (k term) so if you cancel terms you have that S=1/2+1/3- 1/(k+2)- 1/(k+3); you compute the limit when k--> infinity and S=5/6 but cause we factor out 1/2 we have (1/2)*S=(1/2)*(5/6)=5/12

OpenStudy (zzr0ck3r):

there ya go two dif roads lead to same answer:)

OpenStudy (anonymous):

it's not a quite clear explanation but I don't know how to write in LaTex yet ^^

OpenStudy (anonymous):

agree, I think the book is wrong

OpenStudy (zzr0ck3r):

does this all make sense sharazzzz?

OpenStudy (zzr0ck3r):

bah I think hes gone lol

OpenStudy (anonymous):

haha maybe, I hope he'll see this, anyway it's solved

OpenStudy (anonymous):

thanks all of you.. but my question not solve. because i ask the question how to find closed form for a nth partial sum of the series..?

OpenStudy (anonymous):

dear 1/2 k say aaay..?

OpenStudy (zzr0ck3r):

? I have told you the answer 13 times